OpenStudy (anonymous):

Use the pythagorean theorem to determine the third side of the triangle and then find the other five trig functions of sin θ sin θ =5/6

OpenStudy (anonymous):

third side is \[\sqrt{6^2-5^2}=\sqrt{36-25}=\sqrt{11}\]

OpenStudy (anonymous):

both triangles above unfortunately are not right

OpenStudy (anonymous):